OpenStudy (anonymous):

Find two consecutive integers whose sum is 113.

OpenStudy (anonymous):

Find two consecutive integers whose sum is 113.

OpenStudy (anonymous):

let x be the first integer so the consecutive integer would be x + 1 sum is 113 so x + (x+1) = 113 go find x...then find x+1 and you have your numbers

OpenStudy (anonymous):

How far does a plane travel in 3 hours at 654 miles per hour?

OpenStudy (anonymous):

distance is speed times time

OpenStudy (austinl):

654*3=?

OpenStudy (anonymous):

Jude has seven times as many dimes as nickels. Their value is $5.25. How many nickels does he have?

OpenStudy (austinl):

7(0.1x)+(0.05y)=5.25 I believe this is correct. If not someone feel free to correct me.

OpenStudy (anonymous):

Describe the location of the point that has the following coordinates. negative abscissa, zero ordinate

OpenStudy (anonymous):

What is the range of the function y = -|x|?